### v1

# Observation of Self-Bound Droplets of Ultracold Dipolar Molecules

Abstract
Ultracold gases of dipolar molecules have long been envisioned as a platform for the realization of novel quantum phases Góral et al. (2002); Micheli et al. (2006); Büchler et al. (2007, 2007); Cooper and Shlyapnikov (2009); Capogrosso-Sansone et al. (2010); Gorshkov et al. (2011); Syzranov et al. (2014); Schmidt et al. (2022). Recent advances in collisional shielding Gorshkov et al. (2008); Karman and Hutson (2018); Matsuda et al. (2020); Anderegg et al. (2021), protecting molecules from inelastic losses, have enabled the creation of degenerate Fermi gases De Marco et al. (2019); Valtolina et al. (2020); Schindewolf et al. (2022) and, more recently, Bose-Einstein condensation of dipolar molecules Bigagli et al. (2024). However, the observation of quantum phases in ultracold molecular gases that are driven by dipole-dipole interactions has so far remained elusive. In this work, we report the formation of self-bound droplets and droplet arrays in an ultracold gas of strongly dipolar sodium–cesium molecules. Starting from a molecular Bose-Einstein condensate (BEC), microwave dressing fields are used to induce dipole-dipole interactions with controllable strength and anisotropy. By varying the speed at which interactions are induced, covering a dynamic range of four orders of magnitude, we prepare droplets under equilibrium and non-equilibrium conditions, observing a transition from robust one-dimensional (1D) arrays to fluctuating two-dimensional (2D) structures. The droplets exhibit densities up to 100 times higher than the initial BEC, reaching the strongly interacting regime, and suggesting the possibility of a quantum-liquid or crystalline state Gorshkov et al. (2008); Ciardi et al. (2025). This work establishes ultracold molecules as a system for the exploration of strongly dipolar quantum matter and opens the door to the realization of self-organized crystal phases Büchler et al. (2007); Gorshkov et al. (2008); Pupillo et al. (2008) and dipolar spin liquids in optical lattices Yao et al. (2018).
preprint: APS/123-QED
I Introduction
The emergence of droplets is ubiquitous in nature, typically arising from a delicate interplay between attractive and repulsive forces, both in classical and quantum many-body systems. Examples include the formation of galaxies Chandrasekhar and Fermi (1953), stars Kennicutt Jr and Evans (2012), water droplets Young (1805); Stillinger (1980), atomic nuclei Gamow (1930); Weizsäcker (1935); Bohr and Wheeler (1939), and superfluid helium droplets Grebenev et al. (1998); Volovik (2003). In ultracold atomic quantum gases, droplet formation has been observed in bosonic mixtures with short-range van der Waals interactions Petrov (2015); Cabrera et al. (2018), as well as in quantum gases of magnetic atoms with long-range dipolar interactions Schmitt et al. (2016); Chomaz et al. (2016). For magnetic atoms, complex droplet structures have been realized, including droplet arrays Kadau et al. (2016) and supersolid phases Tanzi et al. (2019); Böttcher et al. (2019); Chomaz et al. (2019); Norcia et al. (2021). In both platforms, the droplets are stabilized by repulsive forces originating from quantum fluctuations. Comparing the range of interactions to the interparticle spacing, droplets in atomic quantum gases have so far been dilute and weakly interacting.
Ultracold dipolar molecules are emerging as a powerful platform for many-body physics, quantum simulation, and quantum information DeMille (2002); Carr et al. (2009); Baranov et al. (2012); Cornish et al. (2024). Based on the prospect of realizing many-body systems with strong dipolar interactions, theoretical efforts over the past two decades have pointed out their potential to realize novel quantum phases, such as quantum droplets Schmidt et al. (2022), supersolids Lechner et al. (2015); Lu et al. (2015); Hertkorn et al. (2021), dipolar crystals Büchler et al. (2007); Gorshkov et al. (2008) and strongly correlated lattice models Micheli et al. (2006); Capogrosso-Sansone et al. (2010); Gorshkov et al. (2011). On the experimental side, strong inelastic losses have long posed a major obstacle to creating quantum degenerate gases of molecules Ospelkaus et al. (2010); Bause et al. (2023). Recent breakthroughs came with the development of collisional shielding techniques, using static and oscillating electric fields Matsuda et al. (2020); Anderegg et al. (2021); Bigagli et al. (2023); Lin et al. (2023), which have enabled the creation of degenerate Fermi gases Valtolina et al. (2020); Schindewolf et al. (2022) and Bose-Einstein condensation Bigagli et al. (2024) of dipolar molecules in regimes where dipole-dipole interactions remain relatively weak. However, the creation of degenerate molecular gases with strong dipolar interactions has remained elusive, and it has been a critical open question whether they can be sufficiently long-lived to enable the observation of novel quantum phases.
Figure 1: Droplet formation in an ultracold gas of NaCs molecules with non-axially symmetric dipolar interactions. a, A quantum degenerate gas of microwave-dressed NaCs molecules is confined in an anisotropic optical dipole trap. The molecules are simultaneously dressed with a $\pi$ field with linear polarization along the $z$ direction and a $\sigma$ field with elliptical polarization rotating in the $xy$ plane. The frequencies of the dressing fields are close to 3.5 GHz. By tuning the ellipticity $\xi$ of the $\sigma$ field away from pure $\sigma^{+}$ circular polarization, long-range interactions between the molecules take the form of a dipolar interaction for which the axial symmetry is broken. Larger ellipticity |$\absolutevalue{\xi}$ | leads to stronger dipolar interactions. Gravity points along the $-z$ direction. b, Intermolecular interaction potential in the $xy$ plane, calculated for $\xi=-6^{\circ}$. c, Angular dependence of the short-range part (top) and the long-range part (bottom) of the interaction in three-dimensional space for a fixed intermolecular distance ($\xi=-6^{\circ}$). d, Interaction length scales as a function of $\xi$, calculated from the microwave-dressed interaction potential. The solid (dashed) line shows the dipolar length $a_{\mathrm{d2}}$ (s-wave scattering length $a_{\mathrm{s}}$) as defined in the main text. e, Absorption images of a single droplet (left), a weakly dipolar BEC (middle), and a 1D-droplet array (right) after 25 ms time-of-flight expansion. Each image is an average of three individual shots taken along the $z$ direction. Black ellipses (bottom) indicate the trace of the electric field vector of the $\sigma$ field. Scale bar, 30 µm (e).
Here, we report on the observation of self-bound droplets in an ultracold gas of dipolar sodium–cesium molecules. Starting from a weakly interacting molecular BEC, we use microwave dressing fields to ramp up the interaction strength and observe the formation of both single droplets and droplet arrays. By varying the interaction ramp speed over four orders of magnitude, we observe a transition from stable 1D droplet arrays to fluctuating 2D structures, signaling a crossover from near-adiabatic to non-adiabatic state preparation. In the single-droplet phase, we measure an increase of the molecular density by two orders of magnitude compared to the initial BEC. In this regime, the characteristic range of dipolar interactions significantly exceeds the interparticle spacing, indicating the realization of a dipolar many-body system in the strongly interacting regime.
II Strongly Dipolar Interactions
We control the interactions and collisional properties in our molecular samples using two microwave fields that dress the molecules’ first rotational transitions Karman et al. (2025); Yuan et al. (2025), as illustrated in Fig. 1. The two microwave fields are elliptically-polarized rotating in the $xy$ plane ($\sigma$ field) and linearly-polarized oscillating along the $z$ axis ($\pi$ field), respectively, see Fig. 1a. As the molecular dipoles follow the oscillating electric fields, a time-averaged interaction potential arises that can be well approximated by the sum of a repulsive van der Waals ($\propto 1/r^{6}$) and a dipolar long-range ($\propto 1/r^{3}$) term Karman et al. (2025), see Fig. 1b:
$U_{\mathrm{int}}(\vec{r})=\frac{A(\theta,\varphi)}{r^{\mathrm{6}}}+\frac{B(\theta,\varphi)}{r^{\mathrm{3}}}.$
Here, $r$, $\theta$, and $\phi$ define the relative position vector between two molecules in spherical coordinates, as shown in Fig. 1c.
The van der Waals contribution constitutes a strongly repulsive hardcore potential for intermolecular distances $r\lesssim 2,000\ a_{0}$. The $A(\theta,\varphi)$ term is positive for all angles $\theta$ and $\varphi$ and nearly isotropic (see Methods). The hardcore repulsion shields the molecules against fast inelastic losses and ensures the collisional stability of the molecular gas Gorshkov et al. (2008); Karman and Hutson (2018). In addition, the hardcore potential sets an upper limit for the density that can be reached in the molecular sample.
Figure 2: Single droplets and droplet arrays. After ramping the $\sigma$ field ellipticity to the target value $\xi$ within 200 ms, absorption images after 15, 20, and 25 ms time of flight are shown in the top, middle, and bottom row, respectively. Each image is an average of 3 individual shots, taken along the $z$ direction. The molecular clouds do not show a discernible thermal component due to the high condensate fraction. Below each column, the respective values of $a_{\mathrm{d2}}$ and $a_{\mathrm{s}}$ are shown. (bottom) Illustration of the orientation of the non-axially symmetric dipolar interaction potential with respect to the trap axes and the inferred in-trap density profile of the molecular cloud (gray shading). Scale bar, 30 µm.
The strength and angular dependence of the dipolar contribution is controlled by the microwave parameters as follows: For a perfectly circularly polarized $\sigma$ field (ellipticity $\xi=0$), the strength of the $\pi$ field can be tuned such that dipolar interactions vanish at long range ($B=0$) as the dipolar interactions induced by the $\sigma$ and $\pi$ fields compensate each other Yuan et al. (2025). Under these conditions, where the intermolecular potential is dominated by the hardcore van der Waals potential, we can efficiently cool molecular samples into a BEC, as described in Ref. Bigagli et al. (2024). While keeping the $\pi$ field fixed at this compensation point, changing the ellipticity of the $\sigma$ field induces a dipolar interaction in which the axial symmetry is broken Syzranov et al. (2014); its anisotropy is illustrated in Fig. 1c and described by Chen et al. (2023)
$B(\theta,\varphi)=a_{\mathrm{d2}}(\xi)\frac{\hbar^{2}}{m}{\mathrm{\sqrt{3}\cos 2\varphi\sin^{2}\theta}},$
where $m$ is the mass of the molecule and $\hbar$ the reduced Planck constant. The strength of the dipolar interactions is parametrized by the dipolar length, $a_{\mathrm{d2}}(\xi)\propto\sin 2\xi$. Larger ellipticity leads to stronger dipolar interactions. The sign of $\xi$ encodes the orientation of the polarization ellipse, with positive (negative) values indicating alignment along the $y$ ($x$) axis. The angular anisotropy of $B(\theta,\varphi)$ corresponds to the spherical harmonic $d_{x^{2}-y^{2}}$, which has no axial symmetry along any axis. This form of dipolar interactions is distinctly different from the conventional $d_{3z^{2}-r^{2}}$ form, which is axially symmetric about the $z$ axis and widely used in ultracold gases of magnetic atoms Chomaz et al. (2022). In addition to tuning the ellipticity, there is further flexibility in controlling the anisotropy by tuning the $\pi$ field away from the compensation point, allowing the continuous tuning between axially symmetric and non-axially symmetric dipolar interactions.
[…]
Figure 3: Droplet arrays for slow and fast interaction ramps. a, Experimental sequence. The ellipticity is linearly increased to $5.7^{\circ}$ while the ramp time $t_{\mathrm{ramp}}$ and the associated ramp rate $\dot{\xi}$ are varied. Absorption images of the resulting droplet arrays are recorded after 25 ms time of flight (TOF). b, For various ramp rates $\dot{\xi}$, typical single-shot images (top), averages of 30 individual shots (middle), and averages of the modulus of their Fourier transforms (bottom) are shown. c, Fraction of images showing 2D arrays for a given ramp rate $\dot{\xi}$. Each data point is obtained by counting the fraction of shots (out of a sample of 30 shots) that exhibit 2D arrays. Error bars account for shots for which the array dimensionality is ambiguous. Scale bars, 30 µm for absorption images and 0.1 µm-1 for Fourier transforms (b).
III Expansion dynamics
We distinguish the different phases of the molecular gas by analyzing its expansion dynamics after release from the optical dipole trap. Fig. 2 shows absorption images for various ellipticities $\xi$ after 15, 20, and 25 ms time of flight. While for each of the shown ellipticities there are noticeable differences in the expansion behavior, we phenomenologically group the observed states into three distinct phases:
For ellipticities between $-3.5(5)^{\circ}$ and $3.5(5)^{\circ}$, the expansion dynamics are consistent with a weakly dipolar BEC Stuhler et al. (2005). At $\xi=0(1)\degree$, where dipolar interactions are minimized, the gas expands into an approximately round cloud, as expected for a non-dipolar, repulsively interacting BEC for our trap parameters at the specific expansion times Ketterle et al. (1999). As $\xi$ deviates from zero, we observe electrostriction: the clouds show a significant elongation along the axis of attractive dipolar interactions.
At stronger dipolar interactions, $|\xi|>3.5^{\circ}$, we observe the formation of molecular droplets. The behavior differs markedly for the two orientations of the microwave polarization ellipse, as a result of the interplay between the alignment of the dipolar interaction pattern and the trap anisotropy. For positive (negative) ellipticity the microwave polarization ellipse is aligned along the tighter (looser) trap direction in the $xy$ plane.
For negative ellipticities, $\xi<-3.5^{\circ}$, the gas forms a single self-bound droplet. Compared to the weakly dipolar BEC, the cloud is significantly contracted and shows no observable expansion as a function of time of flight. Taking the finite imaging resolution (see Methods) into account, we infer that the expansion energy is at least 20 times smaller than for the weakly dipolar BEC. Additional images taken along the $x$ axis, see Fig. Extended Data 1, confirm that the droplet is self-bound in all directions. For positive ellipticities, $\xi>3.5^{\circ}$, the gas initially develops a density modulation, which evolves into an array of self-bound droplets at higher ellipticities. During time of flight, the individual droplets do not expand, but move apart from each other due to repulsive side-by-side dipolar interactions Ferrier-Barbut et al. (2016). A quantitative analysis of the cloud sizes during time-of-flight expansion, shown in Fig. Extended Data 2, corroborates the self-bound character of both single-droplet and droplet-array phases.
IV Adiabaticity of interaction ramp
[…]
Next, we investigate the formation of droplet arrays as a function of interaction ramp speed, assessing to which extent the emergence of arrays is a ground-state phenomenon Bisset et al. (2016); Ferrier-Barbut et al. (2018). As shown in Fig. 3a, the ellipticity is ramped linearly from $0^{\circ}$ to $5.7^{\circ}$. With ramp times between $100$ µs and 1 s, the ramp rates $\dot{\xi}$ cover a dynamic range of four orders of magnitude. Fig. 3b shows typical single-shot absorption images after 25 ms time of flight, the average of several individual shots, and the averaged modulus of their Fourier transforms. The latter reveals the spectrum of spatial frequencies and the repeatability of the observed array structures. A comprehensive set of raw images is shown in Fig. Extended Data 3.
For slow ramps, $\dot{\xi}<0.1\degree$/ms, we consistently observe 1D arrays with two to four droplets aligned along the $x$ axis. The spacing between the droplets is stable from shot to shot, as evidenced by the modulation in the averaged Fourier transforms. To confirm the near-equilibrium character of this state, we perform direct evaporative cooling to quantum degeneracy directly at $\xi=+5.7^{\circ}$. We observe arrays that closely match those obtained for slow ramps, suggesting that the 1D droplet array corresponds to the ground state or a low-lying metastable state of the interacting many-body system.
For faster ramps, $\dot{\xi}>0.8\degree$/ms, we observe the emergence of 2D droplet arrays. The number and positions of droplets show strong shot-to-shot fluctuations. In the averaged images the density modulations wash out, reflecting the random character of the fluctuations. This is further evidenced by the absence of higher spatial frequency peaks in the averaged Fourier transforms. Notably, the observed density modulations of the 2D arrays occur both along the directions of repulsive and attractive dipolar interactions, suggesting a mechanism distinct from the roton instability discussed in quantum gases of magnetic atoms Chomaz et al. (2022).
To identify the crossover between near-adiabatic and non-adiabatic preparation, we determine the fraction of shots showing 2D arrays for each $\dot{\xi}$. As shown in Fig. 3c, the transition occurs between $0.12\degree$/ms and $0.41\degree$/ms, corresponding to a ramp time of $t_{\mathrm{ramp}}\sim 20$ ms. This is also confirmed by an analysis of the aspect ratio of the clouds in the averaged images, see Fig. Extended Data 4. These results confirm that the ramp rates used in Fig. 2 ($\dot{\xi}<0.04\degree$/ms) are at least near-adiabatic.
V Droplet Density
Given the strong contraction of the molecular clouds in the droplet phase, it is of particular interest to measure their density (see Fig. 4). Direct in situ imaging is not feasible, as the droplet size is below our imaging resolution. Instead, we leverage the fast switchability of the microwave fields and utilize the density-dependence of two-body loss ($\propto n$) to infer the molecular density. After preparing the sample at a given ellipticity $\xi$, we abruptly turn off the dressing fields within 10 µs (see Fig. 4a), initiating rapid two-body loss of the unshielded molecules Ospelkaus et al. (2010). The molecular decay dynamics is observed for up to 3 ms; Fig. 4b shows examples for the BEC, droplet, and droplet array phases. In both droplet regimes, the decay is significantly faster than in the weakly dipolar BEC. Assuming that many-body correlations do not significantly modify the loss dynamics, enhanced decay corresponds to a higher molecule density Chomaz et al. (2016); Tanzi et al. (2019). Fitting the observed decay curves using a numerical model based on a local density approximation (see Methods), we obtain the mean density $\langle n\rangle$ for each many-body state, as shown in Fig. 4c.
In the weakly dipolar BEC phase, |$\absolutevalue{\xi}$ | < $3.5\degree$, we measure a mean density of $10^{12}\,\mathrm{cm}^{-3}$, consistent with previous results Bigagli et al. (2024). For $\xi<-3.5\degree$, the density increases sharply, possibly indicating a phase transition between the weakly dipolar BEC and the single droplet phase. For $\xi>3.5\degree$, the density also rises as the system enters the droplet array phase, but the change is more gradual. In the single droplet phase, we measure mean densities up to $\sim 10^{14}\,\mathrm{cm}^{-3}$, an increase by two orders of magnitude compared to the BEC, reaching values that are comparable to densities in ultracold atomic gases. At $\xi=-7.1^{\circ}$, the interparticle spacing is $\sim 4,000\,a_{0}$, approaching the lower bound set by the radius of the repulsive hardcore potential, $\sim 2,000\,a_{0}$. This raises the intriguing question whether the droplet approaches a quantum-liquid or crystalline state Gorshkov et al. (2008); Ciardi et al. (2025). The comparison of the dipolar length, $\sim 13,000\,a_{0}$, with the interparticle spacing yields a ratio of 3.6, indicative of a system in the strongly interacting regime.
Finally, we measure the lifetimes for all observed phases (see Fig. Extended Data 5). Even for the densest and most strongly interacting states, the lifetimes exceed 100 ms, which is more than two orders of magnitude longer than the timescale of dipolar interactions, $\sim 1$ ms. This is a remarkable difference to strongly interacting atomic BECs where a large s-wave scattering length in the vicinity of a Feshbach resonance leads to fast recombination losses, limiting the lifetime to less than 1 ms Makotyn et al. (2014). In microwave-dressed molecular samples recombination losses are absent Yuan et al. (2025), highlighting their potential for realizing strongly interacting quantum phases under equilibrium conditions.
VI Conclusions
In conclusion, we have observed self-bound droplets and droplet arrays in a strongly dipolar quantum gas of ultracold molecules. Beyond the expansion dynamics and mean densities discussed here, further investigation of the observed phases will be of great interest. For the strongly interacting single droplet, measurements of vorticity, collective excitations, or shear response could clarify whether the phase has quantum-liquid or crystalline properties Gorshkov et al. (2008); Schmidt et al. (2022); Zhang et al. (2025a); Ciardi et al. (2025). For the droplet arrays, probing phase coherence will elucidate whether the system can show supersolid behavior Tanzi et al. (2019); Böttcher et al. (2019); Chomaz et al. (2019); Norcia et al. (2021); Zhang et al. (2025b). Additional experimental and theoretical studies will be needed to understand the nature of the transitions between the observed phases and the stabilization mechanism of strongly dipolar droplets of microwave-dressed molecules Langen et al. (2025).
This work significantly advances ultracold dipolar molecules as a platform for many-body quantum physics, quantum simulation, and the exploration of self-organization processes in the quantum regime. Microwave dressing provides highly tunable, long-range dipole-dipole interactions, in addition to a hardcore potential which ensures that even dense molecular samples are long-lived. The fast control of microwave fields further opens up new measurement modalities including quenches, non-equilibrium dynamics, and Floquet engineering of interactions. Beyond the 3D bulk systems explored here, exotic forms of self-organization are also expected in lower dimensions Wang et al. (2006); Büchler et al. (2007); Tang et al. (2018) and lattice geometries Micheli et al. (2006); Capogrosso-Sansone et al. (2010); Yao et al. (2018). Our system bridges a gap between well-understood model systems, such as ultracold atoms with contact s-wave interactions, and strongly interacting dense systems, such as superfluid helium Dalfovo and Stringari (2001), electron gases Tsui et al. (1982), and nuclear matter Hammer et al. (2013).
VII Acknowledgements
We acknowledge helpful discussions with Blair Blakie, Francesca Ferlaino, Tim Langen, Andrew Millis, Tilman Pfau, Wilhelm Zwerger, and Martin Zwierlein. We thank Lin Su and Tarik Yefsah for critical reading of the manuscript. We are grateful to Aden Lam and Claire Warner for important contributions in the construction of the experimental apparatus. This work was supported by an NSF CAREER Award (Award No. 1848466), an NSF Single Investigator Award (Award No. 2409747), an ONR DURIP Award (Award No. N00014-21-1-2721), an AFOSR Single Investigator Award (Award No. FA9550-25-1-0048), and a grant from the Gordon and Betty Moore Foundation (Award No. GBMF12340). I.S. was supported by the Ernest Kempton Adams Fund. T.K. acknowledges NWO VIDI (Grant ID 10.61686/AKJWK33335). S.W. acknowledges additional support from the Alfred P. Sloan Foundation.